A 42.1 g piece of metal alloy absorbs 181 J as its temperature increases from 25.0 C to 36.0 C. What is the alloy’s specific heat?


1
Expert's answer
2022-04-27T09:37:03-0400

36-25=11 C=11 K

Alloy's specific heat = Q/(mT)=181/(11x42.1)=0.39 J/(gK)